### v1

@@
Linghong Lin 11    Beibei Liu 11⋆⋆    Zekai Zheng 1122
We present a theoretical model to investigate a two-planet pair that undergoes convergent type I migration in a gaseous protoplanetary disk and traps into the first-order mean motion resonance. Our study identifies the conditions for resonant capture and explores the subsequent dynamical stability of the system. We derive the analytical criteria for planets with an arbitrary mass ratio and validate through numerical N-body simulations. Slow migration and weak eccentricity damping are required for resonant capture, the latter of which has not received enough attention in the literature. Once capture into resonance, their following stability can be classified into three regimes: stable trap, overstable trap and escape. Notably, resonant capture remains stable when the inner planet significantly outweighs the outer one. However, the subsequent evolution can be diverse when the mass of the inner planet is lower or comparable to that of the outer planet. Stability weakens as the relative strength between migration and eccentricity damping increases. The key results can be comprehensively demonstrated in a $\tau_{a}$-$\tau_{a}/\tau_{\rm e}$ plot, where $\tau_{a}$ and $\tau_{\rm e}$ are orbital decay and eccentricity damping timescales, respectively.
@@ 1 Introduction
Low-mass planets embedded in gaseous protoplanetary disks interact with disk gas and undergo type I orbital migration (Kley & Nelson, 2012; Paardekooper et al., 2023). Resonance capture can be a natural outcome of the planet pairs featured with convergent migration (Lee & Peale, 2002). The key question is to understand the under which conditions this capture can occur. The classical pendulum model pointed out that the trapping holds when the time planet migrates across the resonance width is longer than the time resonant angle librates (Murray & Dermott, 1999; Ogihara & Kobayashi, 2013; Batygin, 2015; Liu et al., 2015). A new criterion was derived by Huang & Ormel (2023), where they analyzed the resonant equilibrium state under the disk migration effects in a restrict three-body framework. Batygin & Petit (2023) provided an equivalent criterion using the Hamiltonian theory and extended it to the context of the unrestricted problem.
@@
On the other hand, Goldreich & Schlichting (2014) first investigated the long-term stability of the resonant pair after capture in a gaseous disk. They found that under certain conditions, the resonant equilibrium state can become overstable, indicating that the planets may temporarily or permanently escape from resonance. Deck & Batygin (2015) conducted a similar analysis based on the Hamiltonian theory. Xu & Lai (2017) extended the framework to second-order mean motion resonances. Their study highlighted that the mass ratio of the planet pair is a crucial parameter in determining the overstability. Nesvorný et al. (2022) applied the overstable libration to account for the high eccentricity and large amplitude of the resonant angle observed in TOI-216b. Recently, Afkanpour et al. (2024) verified this overstable resonant trap using hydrodynamical simulations. They found that the outcome is related to the disk properties, such as disk mass and viscosity.
@@
In this paper, we develop a unified framework for a two-planet pair that undergoes convergent type I migration and gets trapped at the first-order mean motion resonance ($j$:$j-1$ MMR). We investigate the conditions under which planets can be captured in resonances and examine the following stability properties. The key analytical criteria are derived for planets with arbitrary mass ratios and the outcomes are summarized in a $\tau_{\rm a}{-}\tau_{\rm a}/\tau_{\rm e}$ diagram.
@@
The paper is organized as follows. We present an analytical model accounting for the dynamics of planets at resonances in Sec. 2. In Sec. 3 we perform N-body simulations and compare the results with the analytical predictions. In Sec. 4 we discuss several implications as well as potential caveats in our model. Finally, we summarize the key results in Sec. 5.
2 Dynamics of resonance trapping
@@
In this section, we discuss the dynamical evolution of a two-planet pair in a gaseous protoplanetary disk after their convergent type-I migration. We consider two planets of masses $m_{\rm i}$ and $m_{\rm o}$ orbiting a central star of $M_{\star}$ on coplanar orbits. Their reduced masses and the mass ratio between the inner and outer ones are $\mu_{\rm i,o}{=}m_{\rm i,o}/M_{\rm\star}$ and $q{=}m_{\rm i}/m_{\rm o}$, respectively. For notations the subscripts ‘i’ and ‘o’ refer to the corresponding quantities of inner and outer planets.
@@
We present the analytical criteria of resonance trapping and stability in Sec. 2.1 and Sec. 2.2, respectively. A unified picture is summarized in Sec. 2.3.
@@ 2.1 Resonance trapping criteria
The dissipative effect of planet-disk interaction can be parameterized by (Teyssandier & Terquem, 2014)
$\frac{1}{a}\frac{da}{dt}$ $=$ $-\frac{1}{\tau_{a}}-\frac{2e^{2}}{\tau_{e}},$ (1a)
$\frac{1}{e}\frac{de}{dt}$ $=$ $-\frac{1}{\tau_{e}},$ (1b)
@@
where $\tau_{a}$ and $\tau_{e}$ are the orbital decay and eccentricity damping timescales, respectively. The orbital elements $a$, $n$, $e$, $\lambda$, $\varpi$, denote the semimajor axis, mean motion, eccentricity, mean longitude and longitude of pericenter of the planet.
@@
Figure 1: Resonance crossing of a two-planet system with convergent migration. The plots provide the time evolution of the outer-to-inner planet period ratio (panel a), the two planets’ eccentricities (panel b), resonant angle (panel c) and phase curve (panel d). The initial period ratio of the two planets is $2.2$, just outside of the 2:1 MMR, the planet masses are $m_{\rm i}=1\ M_{\oplus}$, $m_{\rm o}=10\ M_{\oplus}$, and the migration parameters are $\tau_{a}=1.1\times 10^{5}yr$ and $\tau_{a}/\tau_{e}=1.5\times 10^{3}$. The planets directly cross the $2$:$1$ resonance at $t{\simeq}6\times 10^{3}$ yr and their eccentricities simultaneously excite to $10^{-2}$ and rapidly decline afterward. The resonant angle of the inner planet jumps from the initial values to $\pi$ during the crossing and oscillates over $2\pi$ afterwards. In phase space, the inner planet is not captured by the non-zero resonant fixed point and ultimately evolves towards the origin.
@@
For the $j$:$j-1$ order mean motion resonances (MMRs), the resonant angles and the time-averaged disturbing functions for the inner and outer planets are given by (Murray & Dermott, 1999) $\varphi_{\rm i}{=}j\lambda_{\rm o}+(1-j)\lambda_{\rm i}-\varpi_{\rm i}$, $\varphi_{\rm o}{=}j\lambda_{\rm o}+(1-j)\lambda_{\rm i}-\varpi_{\rm o}$, and $R_{\rm i}{=}Gm_{\rm o}/a_{\rm o}(R_{\rm D}^{\rm sec}+R_{\rm D}^{\rm res}+\alpha R_{\rm E})$, $R_{\rm o}{=}Gm_{\rm i}/a_{\rm o}(R_{\rm D}^{\rm sec}+R_{\rm D}^{\rm res}+R_{\rm I}/\alpha^{2})$, where $\alpha{=}a_{\rm i}/a_{\rm o}$ is the two planets’ semi-major axis ratio, $R_{\rm D}^{\rm sec}$ and $R_{\rm D}^{\rm res}$ are the secular and resonant terms, $R_{\rm E}$ and $R_{\rm I}$ are the indirect part due to an external and internal perturber.
@@
We assume small eccentricities of the planets, and expand the disturbing functions to the first order in $e_{\rm i}$ and $e_{\rm o}$, which gives $R_{\rm D}^{\rm sec}{=}0$, $R_{\rm D}^{\rm res}{=}e_{\rm i}f_{\rm d,i}\cos{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}\cos{\varphi_{\rm o}}$,$R_{\rm E}{=}-2e_{\rm o}\cos{\varphi_{\rm o}}\delta_{\rm j,2}$, and $R_{\rm I}{=}-\frac{1}{2}e_{\rm o}\cos{\varphi_{\rm o}}\delta_{\rm j,2}$, where $\delta_{\rm j,2}$ is the usual Kronecker symbol. Detailed expressions of $f_{\rm d,i}$ and $f_{\rm d,o}$ can be found in Table 8.1 of Murray & Dermott (1999).
@@
$\dot{n_{\rm i}}$ $=-3(j-1)\mu_{\rm o}n_{\rm i}^{2}\alpha(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})+\frac{3n_{\rm i}}{2\tau_{a,\rm i}}+\frac{3n_{\rm i}e_{\rm i}^{2}}{\tau_{e,\rm i}},$ (2a)
$\dot{n_{\rm o}}$ $=3j\mu_{\rm i}n_{\rm o}^{2}(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})+\frac{3n_{\rm o}}{2\tau_{a,\rm o}}+\frac{3n_{\rm o}e_{\rm o}^{2}}{\tau_{e,\rm o}},$ (2b)
@@
where $f_{\rm d,o}^{{}^{\prime}}{=}f_{\rm d,o}-2\alpha\delta_{j,2}$. For instance, $f_{\rm d,i}{=}{-}1.19$, $f_{\rm d,o}^{{}^{\prime}}{=}0.43$ at the $2$:$1$ MMR.
When the planets are in MMRs, their orbital parameters reach the equilibrium state such that $\dot{n_{\rm i}}/n_{\rm i}{=}\dot{n_{\rm o}}/n_{\rm o}$, $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$, $\dot{\varpi_{\rm i}}{=}\dot{\varpi_{\rm o}}$. By substituting $\dot{n_{\rm i}}/n_{\rm i}{=}\dot{n_{\rm o}}/n_{\rm o}$ into Laplace equation, Eqs. (2a) and (2b) can be combined as
@@
$=\frac{1}{2\tau_{a}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}},$
where we assume that $\tau_{a}^{-1}{=}\tau_{a,\rm o}^{-1}-\tau_{a,\rm i}^{-1}$ is the relative orbital decay rate with respect to convergent migration, and $\tau_{e}{=}\tau_{e,\rm i}{=}\tau_{e,\rm o}$ represents the eccentricity damping time-scale. Substituting $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$ into Eqs. (2c) and (2d) to eliminate $\sin{\varphi_{\rm i,o}}$ and $\dot{\varpi_{\rm i}}{=}\dot{\varpi_{\rm o}}$ into Eqs. (2e) and (2f) to eliminate either $e_{\rm i}$ or $e_{\rm o}$, the equilibrium eccentricities can be derived from Eq. (3) such that
$e_{\rm i,eq}^{2}$ $=$ $\frac{\tau_{e}}{2\tau_{a}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (4a)
$e_{\rm o,eq}^{2}$ $=$ $\frac{\tau_{e}}{2\tau_{a}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (4b)
@@
where $e_{\rm o,eq}/e_{\rm i,eq}{=}{q\sqrt{\alpha}|f_{\rm d,o}^{{}^{\prime}}/f_{\rm d,i}|}$.
@@
Our results are analogous to Terquem & Papaloizou (2019), except for $\tau_{a}$ and $\tau_{e}$ as defined. In the limit of the mass of inner planet much lower than the outer one ($q{\ll}1$), the eccentricity of inner planet in Eq. (4a) can be simplified as
$e_{\rm i,eq}=\sqrt{\frac{1}{2j}\frac{\tau_{e}}{\tau_{a}}},$ (5)
@@
in agreement with Eq. (24) in Huang & Ormel (2023). We note that due to differing signs of $f_{\rm d,i}$ and $f_{\rm d,o}^{{}^{\prime}}$, $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$ requires that $\sin{\varphi_{\rm i}}{>}0$ and $\sin{\varphi_{\rm o}}{<}0$ for $j$:$j-1$ MMRs. Considering $e_{\rm o,eq}/e_{\rm i,eq}{=}{q\sqrt{\alpha}|f_{\rm d,o}^{{}^{\prime}}/f_{\rm d,i}|}$, we can derive $\sin{\varphi_{\rm i}}{=}-\sin{\varphi_{\rm o}}$, which implies $\varphi_{\rm i}-\varphi_{\rm o}{=}\pm\pi$. This means that the conjunctions between planet pair in a mean motion resonance always occur at either pericentre or apocentre.
@@
Notably, the right-hand side of Eq. (3) represents the planet-disk interaction that leads convergent orbital migration between two planets, while the left-hand side of Eq. (3) represents their resonant gravitational interaction, keeping them away from getting closer. When the effect of planet-disk interaction is strong enough to overcome the resonant interaction, the planet pair cannot stay in resonance. Strong planet-disk interaction can be further classified into two regimes: fast migration and strong eccentricity damping. In the first regime the exchange rate of semimajor axis $\dot{a}/a$ in Eq. (3) is driven by the a-decay term ($\tau_{a}^{-1}{\gg}e^{2}\tau_{e}^{-1}$). In the second regime, however, the contribution of the e-damping term becomes comparable to that of the a-decay term. We will derive the resonance trapping conditions in the following subsections.
Figure 2: Resonance crossing of a two-planet system with convergent migration. The initial period ratio of the two planets is $2.2$, just outside of the 2:1 MMR, and the migration parameters are $\tau_{a}=10^{5}yr$ and $\tau_{a}/\tau_{e}=6\times 10^{2}$. The planets are trapped in the $2:1$ resonance at $t{\simeq}10^{4}$ yr and their eccentricities simultaneously excite to $2\times 10^{-2}$, which remain stable thereafter. The resonant angle of the inner planet is excited when the period ratio evolves to 2:1, subsequently oscillating and gradually converging to the resonant equilibrium value, which it maintains. In phase space, the inner planet is captured by the non-zero resonant fixed point and subsequently evolves in conjunction with this fixed point, reflecting the static equilibrium state of the resonance..

### v2

@@
Linghong Lin 11    Beibei Liu 1 ⋆1 ⋆    Zekai Zheng 1122
We present a theoretical framework for investigating a two-planet system undergoing convergent type I migration in a protoplanetary disk. Our study identifies the conditions for resonant capture and subsequent dynamical stability. By deriving analytical criteria for general $j$:j1j-1 first-order mean-motion resonances (MMRs) applicable to planet pairs with arbitrary mass ratios, we validate these predictions through N-body simulations. The key results are demonstrated in $\tau_{\rm m}$-$\tau_{\rm m}/\tau_{e}$ plots, where $\tau_{\rm m}$ and $\tau_{e}$ are the timescales of the angular momentum and eccentricity damping, respectively. Specifically, we determine which combinations of orbital damping timescales allow for capture into resonance, showing that too fast migration or too strong eccentricity damping inhibit successful capture. After capture, the subsequent evolution can be classified into three regimes: stable trap, overstable trap and escape. Importantly, resonant capture always remains stable when the inner planet significantly outweighs the outer one. In contrast, when the mass of the inner planet is lower than or comparable to that of the outer planet, the system transitions from the stable to overstable trap, and eventually escapes the resonance, as the relative strength of eccentricity damping to migration ($\tau_{\rm m}/\tau_{e}$) decreases.
@@ 1 Introduction
Low-mass planets embedded in protoplanetary disks interact with surrounding disk gas and undergo type I orbital migration (Goldreich & Tremaine, 1980; Ward, 1997; Tanaka et al., 2002; Kley & Nelson, 2012; Baruteau et al., 2014; Paardekooper et al., 2023). Resonance capture can be a natural outcome of a planet pair that undergoes convergent migration (Lee & Peale, 2002; Papaloizou & Szuszkiewicz, 2005). The key question is to understand under which conditions this capture can occur and its following dynamical evolution.
@@
The classical pendulum model intuitively shows that the trapping holds when the time planet migrates across the resonance width is longer than the libration period in the resonance (Murray & Dermott, 1999; Ogihara & Kobayashi, 2013; Liu et al., 2015), which is akin to the adiabatic condition described by Batygin (2015). The eccentricity damping from the gas disk is neglected in these studies. Later, Kajtazi et al. (2023) demonstrated through simulations that eccentricity damping qualitatively alters the capture conditions. Incorporating eccentricity damping, Huang & Ormel (2023) derived a new criterion by analyzing the resonant equilibrium state under disk-driven migration within a restricted three-body framework. Meanwhile, Batygin & Petit (2023) derived an equivalent criterion using Hamiltonian theory and extended the analysis to the two-planet pairs with arbitrary mass ratios.
@@
On the other hand, Goldreich & Schlichting (2014) first investigated the long-term stability of the resonant planet pair after capture in a gaseous disk. Their study revealed that under certain conditions, the resonant equilibrium state can become overstable, where the planets may eventually escape from resonance. Building on this, Deck & Batygin (2015) conducted a similar analysis using Hamiltonian theory and extended the work to the unrestricted three-body system. Xu & Lai (2017) further expanded the framework to second-order mean motion resonances. These studies highlighted that the mass ratio of the planet pair is a crucial parameter in determining whether the system is overstable. Nesvorný et al. (2022) applied the overstable libration to account for the high eccentricity and large amplitude of the resonant angle observed in TOI-216b. Recently, Afkanpour et al. (2024) verified this overstable resonant trap using hydrodynamical simulations. They found that the outcome is related to disk properties, such as the gas mass and turbulent viscosity.
@@
In this paper, we develop a unified framework for a two-planet pair that undergoes convergent type I migration and gets trapped at the first-order mean motion resonance ($j$:j1j-1 MMR). We investigate the conditions under which planets can be captured in resonances and examine the subsequent stability properties. The key analytical criteria are derived for planets with arbitrary mass ratios, and the outcomes are summarized in a $\tau_{\rm m}{-}\tau_{\rm m}/\tau_{e}$ parameter space.
@@
The paper is organized as follows. In Section 2, we present an analytical model that describes the dynamics of planets in resonant configurations. In Section 3 we perform N-body simulations and compare the results with analytical predictions. Section 4 outlines potential caveats, and Section 5 summarizes our key results.
2 Dynamics of resonance capture and stability
@@
This section discusses the dynamical evolution of a two-planet pair in a gaseous disk through the convergent type-I migration. We consider two planets of masses $m_{\rm i}$ and $m_{\rm o}$ orbiting a central star of $M_{\star}$ on coplanar orbits. Their normalized masses and the mass ratio between the inner and outer ones are $\mu_{\rm i,o}{=}m_{\rm i,o}/M_{\rm\star}$ and $q{=}m_{\rm i}/m_{\rm o}$, respectively. For notation, the subscripts ‘i’ and ‘o’ refer to the corresponding quantities of the inner and outer planets.
@@
We present the analytical criteria for resonance trapping and stability in Section 2.1 and Section 2.2, respectively. A unified picture is summarized in Section 2.3. Resonance crossing in the context of divergent migration is also discussed in Section 2.4.
@@ 2.1 Resonance trapping criteria
The dissipative effect of the planet-disk interaction can be parameterized by (Teyssandier & Terquem, 2014; Ataiee & Kley, 2021; Pichierri et al., 2023)
$\frac{1}{L}\frac{dL}{dt}$ $=$ $-\frac{1}{\tau_{\rm m}},$ (1a)
$\frac{1}{a}\frac{da}{dt}$ $=$ $-\frac{1}{\tau_{\rm a}}=-\frac{2}{\tau_{\rm m}}+\frac{2e^{2}}{1-e^{2}}\frac{1}{\tau_{e}},$ (1b)
$\frac{1}{e}\frac{de}{dt}$ $=$ $-\frac{1}{\tau_{e}},$ (1c)
@@
where the orbital elements $L$, $a$, $n$, $e$, $\lambda$, $\varpi$, denote the angular momentum, semimajor axis, mean motion, eccentricity, mean longitude and longitude of pericenter of the planet, $\tau_{\rm m}$, $\tau_{\rm a}$ and $\tau_{\rm e}$ represent the damping timescales for orbital angular momentum, semi-major axis and eccentricity, respectively. It is worth noting that in the limit of small eccentricity, the angular momentum and orbital decay timescales can be linked through $\tau_{\rm m}{\approx}2\tau_{\rm a}$.
@@
Figure 1: Resonance crossing of a two-planet system with convergent migration. The plot shows the time evolution of (a) the outer-to-inner planet period ratio, (b) the planet eccentricities, (c) the resonant angle, and (d) the phase curve. The planet and disk parameters are $m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $\tau_{\rm m}{=}2.2\times 10^{5}$ yr and $\tau_{\rm m}/\tau_{e}{=}3\times 10^{3}$. The planets directly cross the 22:11 resonance at $t{\simeq}6\times 10^{3}$ yr, while the inner planet’s eccentricity briefly excites to $0.01$ before rapid decay. During the crossing, the resonant angle of the inner planet $\varphi_{\rm i}$ jumps from 0 to $\pi$ and subsequently oscillates with growing amplitude22footnotemark: 2. In phase space, the planet’s trajectory is not captured by the resonant equilibrium. Instead, it circulates and eventually evolves toward the origin, corresponding to a low-eccentricity, non-resonant state. See https://github.com/llh-astro/resonance/raw/main/gif/no_trap.gif
@@
For a $j$:j1j-1 MMR, the resonant angles and the time-averaged disturbing functions for the inner and outer planets are given by (Murray & Dermott, 1999) $\varphi_{\rm i}{=}j\lambda_{\rm o}+(1-j)\lambda_{\rm i}-\varpi_{\rm i}$, $\varphi_{\rm o}{=}j\lambda_{\rm o}+(1-j)\lambda_{\rm i}-\varpi_{\rm o}$, and $R_{\rm i}{=}Gm_{\rm o}/a_{\rm o}(R_{\rm D}^{\rm sec}+R_{\rm D}^{\rm res}+\alpha R_{\rm E})$, $R_{\rm o}{=}Gm_{\rm i}/a_{\rm o}(R_{\rm D}^{\rm sec}+R_{\rm D}^{\rm res}+R_{\rm I}/\alpha^{2})$, where $\alpha{=}a_{\rm i}/a_{\rm o}$ is the two planets’ semi-major axis ratio, $R_{\rm D}^{\rm sec}$ and $R_{\rm D}^{\rm res}$ are the secular and resonant terms, $R_{\rm E}$ and $R_{\rm I}$ are the indirect part due to an external and internal perturber.
@@
We assume small eccentricities of the planets, and expand the disturbing functions to the first order in $e_{\rm i}$ and $e_{\rm o}$, which gives $R_{\rm D}^{\rm sec}{=}0$, $R_{\rm D}^{\rm res}{=}e_{\rm i}f_{\rm d,i}\cos{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}\cos{\varphi_{\rm o}}$$R_{\rm E}{=}-2\delta_{\rm j,2}e_{\rm o}\cos{\varphi_{\rm o}}$, and $R_{\rm I}{=}-\frac{1}{2}\delta_{\rm j,2}e_{\rm o}\cos{\varphi_{\rm o}}$, where $\delta_{\rm j,2}$ is the usual Kronecker symbol. Detailed expressions of $f_{\rm d,i}$ and $f_{\rm d,o}$ can be found in Table 8.1 of Murray & Dermott (1999).
@@
$\dot{n_{\rm i}}$ $=-3(j-1)\mu_{\rm o}n_{\rm i}^{2}\alpha(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})+\frac{3n_{\rm i}}{\tau_{\rm m\rm,i}}+\frac{3n_{\rm i}e_{\rm i}^{2}}{\tau_{e,\rm i}},$ (2a)
$\dot{n_{\rm o}}$ $=3j\mu_{\rm i}n_{\rm o}^{2}(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})+\frac{3n_{\rm o}}{\tau_{\rm m\rm,o}}+\frac{3n_{\rm o}e_{\rm o}^{2}}{\tau_{e,\rm o}},$ (2b)
@@
where $f_{\rm d,o}^{{}^{\prime}}{=}f_{\rm d,o}-2\alpha\delta_{j,2}$. For instance, $f_{\rm d,i}{=}{-}1.19$, $f_{\rm d,o}^{{}^{\prime}}{=}0.43$ at the 22:11 MMR.
The orbital parameters of the planets reach the equilibrium state at resonances such that $\dot{n_{\rm i}}/n_{\rm i}{=}\dot{n_{\rm o}}/n_{\rm o}$, $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$, $\dot{\varpi_{\rm i}}{=}\dot{\varpi_{\rm o}}$. By substituting $\dot{n_{\rm i}}/n_{\rm i}{=}\dot{n_{\rm o}}/n_{\rm o}$ into Lagrange’s equation, Eqs. (2a) and (2b) can be combined as
@@
$=\frac{1}{\tau_{\rm m}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}},$
where we define that $\tau_{\rm m}^{-1}{=}\tau_{\rm m,o}^{-1}-\tau_{\rm m,i}^{-1}$ is the relative angular momentum damping rate for convergent migration, and $\tau_{e}=\tau_{e,\rm i}=\tau_{e,\rm o}$ represents the eccentricity damping timescale. In reality, the eccentricity damping rates of the two planets may vary, depending on disk properties and planet masses. For simplicity, we assume equal $\tau_{e}$ for both planets in our theoretical derivations and numerical simulations in the main text. We note, however, that our theoretical framework can be naturally extended to the case of unequal eccentricity damping. The expressions of our criteria for different eccentricity damping timescales are elaborated in Appendix C.
Substituting $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$ into Eqs. (2c) and (2d) to eliminate $\sin{\varphi_{\rm i,o}}$ and $\dot{\varpi_{\rm i}}{=}\dot{\varpi_{\rm o}}$ into Eqs. (2e) and (2f) to eliminate either $e_{\rm i}$ or $e_{\rm o}$. Here we assume $\cos{\varphi_{\rm i}}=\cos{\varphi_{\rm o}}=1$. Under these conditions, the equilibrium eccentricities can be derived from Eq. (3), yielding
$e_{\rm i,eq}^{2}$ $=$ $\frac{\tau_{e}}{\tau_{\rm m}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (4a)
$e_{\rm o,eq}^{2}$ $=$ $\frac{\tau_{e}}{\tau_{\rm m}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (4b)
@@
where $e_{\rm o,eq}/e_{\rm i,eq}{=}{q\sqrt{\alpha}|f_{\rm d,o}^{{}^{\prime}}/f_{\rm d,i}|}$ (see Lee 2004; Deck & Batygin 2015; Ramos et al. 2017; Terquem & Papaloizou 2019).
@@
In the limit of the mass of the inner planet much lower than the outer one ($q{\ll}1$), the eccentricity of the inner planet in Eq. (4a) can be simplified as
$e_{\rm i,eq}=\sqrt{\frac{1}{j}\frac{\tau_{e}}{\tau_{\rm m}}}.$ (5)
@@
We note that due to differing signs of $f_{\rm d,i}$ and $f_{\rm d,o}^{{}^{\prime}}$, $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$ requires that $\sin{\varphi_{\rm i}}{>}0$ and $\sin{\varphi_{\rm o}}{<}0$ for $j$:j1j-1 MMRs. Considering $e_{\rm o,eq}/e_{\rm i,eq}{=}{q\sqrt{\alpha}|f_{\rm d,o}^{{}^{\prime}}/f_{\rm d,i}|}$, we can derive $\sin{\varphi_{\rm i}}{=}-\sin{\varphi_{\rm o}}$, which implies $\varphi_{\rm i}-\varphi_{\rm o}{=}\pm\pi$. This means that the conjunctions between planet pairs in a mean motion resonance always occur at either pericentre or apocentre.
@@
In particular, the right-hand side of Eq. (3) represents the planet-disk interaction that leads to convergent orbital migration between two planets, while the left-hand side of Eq. (3) represents their resonant gravitational interaction. When the effect of planet-disk interaction is strong enough to overcome the resonant interaction, the planet pair cannot stay in resonance. The strong planet-disk interaction can be classified into two regimes: fast migration and strong eccentricity damping. In the first regime, $\dot{a}/a$ in Eq. (3) is driven by the angular momentum damping term ($\tau_{\rm m}^{-1}{\gg}e^{2}\tau_{e}^{-1}$). In the second regime, however, the contribution of the e-damping term becomes comparable to that of the angular momentum damping term. We will derive the resonance trapping conditions from these two regimes in the following subsections.
Figure 2: Stable resonant capture of a two-planet system with convergent migration. The planet and disk parameters are $m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $\tau_{\rm m}{=}2\ \times 10^{5}$ yr and $\tau_{\rm m}/\tau_{e}{=}1200$. The planets get trapped in a 22:11 MMR at $t{\simeq}10^{4}$ yr exciting the inner planet’s eccentricity to $0.02$. The resonant angle $\varphi_{\rm i}$ shifts from 0 to $2\pi/3$ as the system transitions into the resonant state. It gradually decreases as convergent migration strengthens the resonant interaction, causing the libration center to drift slowly. In phase space, the trajectory converges toward the resonant equilibrium and co-evolves with it. We refer to this resonant capture at a stationary equilibrium as a stable trap. See https://github.com/llh-astro/resonance/raw/main/gif/stable_trap.gif